OpenStudy (heart_offire):

Suppose a line has slope -2/3 and passes through the point (1, 8). Which other point must also be on the graph? A) (7, 4) B) (7, 12) C) (-5, 4) D) (-5, -12)

OpenStudy (heart_offire):

@jim_thompson5910

OpenStudy (photon336):

you can use the point slope formula

OpenStudy (heart_offire):

ok but dont i need two sets of points instead of one

OpenStudy (photon336):

not necessarily all you need is the slope of the line and one of the points that you want it to go through \[m(x-x_{0}) = (y-y_{0})\] where m = -2/3 and x0,y0 = (1,8) \[\frac{ -2 }{ 3 }(x-1) = (y-8)\]

OpenStudy (photon336):

now simplify the equation

OpenStudy (heart_offire):

@Photon336 i got B

OpenStudy (photon336):

how did you get that ? also first simply the equation above

OpenStudy (heart_offire):

im kindve confused on simplifying, i turned the fraction into a decimal

satellite73 (satellite73):

slope of \(-\frac{2}{3}\) means right three , down two

satellite73 (satellite73):

so from \((1,8)\) right three down two takes you to \((4,5)\) you can find the others that way

satellite73 (satellite73):

ok actually that was wrong' from \((1,8)\) right three down two goes to \((4,6)\)

satellite73 (satellite73):

go right another three ,down another two where do you end up /

OpenStudy (heart_offire):

umm 7,4

satellite73 (satellite73):

um yeah

OpenStudy (heart_offire):

oh ok so what next

satellite73 (satellite73):

that was all the question asked right? what other point is on the line

OpenStudy (heart_offire):

oh ok

OpenStudy (heart_offire):

thanks